A 2.0 kg ball moving east at 0.5 m/s collides with a 3.0 kg ball moving west at 0.75 m/s. After the collision, the lighter ball moves west at 0.65 m/s. What is the velocity of the larger ball after the collision?

1
Expert's answer
2021-12-14T09:27:23-0500

"m_1v_1-m_2v_2=-m_1v_1'+m_2v_2'\to"


"v_2'=(m_1v_1-m_2v_2+m_1v_1')\/m_2="


"=(2\cdot0.5-3\cdot0.75+2\cdot0.65)/3=0.017\ (m/s)" eastward